Charge q1 = 7.00 μC is at the origin, and charge q2 = -5.00 μC is on the x-axis, 0.300 m from the origin as shown below. (a) Find the magnitude and direction of the electric field at point P, which has coordinates (0, 0.400) m. (b) Find the force on a charge of 2.00 × 10-8 C placed at P


1
Expert's answer
2022-02-02T05:15:22-0500

a)

"E=E_1-E_2=-k(\frac {q_1}{r_1^2}-\frac {q_2}{r_2^2})=4.1~\frac {MN}C,"

b)

"F=q_3E=-0.082~N."